OpenStudy (anonymous):

How do I finish solving this ? log(t−50160)=h⋅log(0.95)

OpenStudy (solomonzelman):

\[a~Log~(c)=Log~(c^a)\]use this on the right side.

OpenStudy (anonymous):

that is actuallly log (t-50/(160))

OpenStudy (solomonzelman):

Ok, do the right side as I told you, what do you get?

OpenStudy (anonymous):

log (0.95^h

OpenStudy (solomonzelman):

YEP!\[Log(0.95^h)~~~~~~~~~~~~~~~~~Is~~~right!\]

OpenStudy (solomonzelman):

\[If~~~~Log(a)=Log(b)~~~~~~~~~~~~~then~~a=b.~~~~~~(Apply~~this~~rule~~here.)\]

OpenStudy (anonymous):

What do you mean ? Like both logs are equal

OpenStudy (solomonzelman):

yes, both numbers inside are equal just like a=b.

OpenStudy (solomonzelman):

\[0.95^h=\frac{t-50}{160}\]

OpenStudy (anonymous):

so the answer is .0167

OpenStudy (solomonzelman):

How are you getting the answer, if you have 2 unknown variables in 1 equation?

OpenStudy (anonymous):

Okay never mind I was taking the log of .95

OpenStudy (solomonzelman):

without times h?

OpenStudy (anonymous):

yes I thought what you meant by equal was that I could do that but I was wrong

OpenStudy (solomonzelman):

Could do what, remove the h?

OpenStudy (solomonzelman):

I didn't mean that...

OpenStudy (solomonzelman):

If you wrote it correctly, then that's the answer.

OpenStudy (anonymous):

Okay so how do I move it to one side. This was me trying to make the inverse of this function h(t)=106(0.95)^t+50

OpenStudy (solomonzelman):

how do you move it to one side? Like what?

OpenStudy (anonymous):

Like for it to be inverse of the original function

OpenStudy (solomonzelman):

OK, inverse each action/step going from the first term (this should make sense if you've done this staff before) Like if it is f(x)=2x+3